### v2

# LeanQuant: Accurate and Scalable Large Language Model Quantization with Loss-error-aware Grid

Abstract
Large language models (LLMs) have shown immense potential across various domains, but their high memory requirements and inference costs remain critical challenges for deployment. Post-training quantization (PTQ) has emerged as a promising technique to reduce memory requirements and decoding latency. However, recent accurate quantization methods often depend on specialized computations or custom data formats to achieve better model quality, which limits their compatibility with popular frameworks, as they require dedicated inference kernels tailored to specific hardware and software platforms, hindering wider adoption. Furthermore, many competitive methods have high resource requirements and computational overhead, making it challenging to scale them to hundreds of billions of parameters. In response to these challenges, we propose LeanQuant (Loss-error-aware Network Quantization), a novel quantization method that is accurate, versatile, and scalable. In the existing popular iterative loss-error-based quantization framework, we identify a critical limitation in prior methods: the min-max affine quantization grid fails to preserve model quality due to outliers in inverse Hessian diagonals. To overcome this fundamental issue, we propose learning loss-error-aware grids, instead of using non-adaptive min-max affine grids. Our approach not only produces quantized models that are more accurate but also generalizes to a wider range of quantization types, including affine and non-uniform quantization, enhancing compatibility with more frameworks. Extensive empirical evaluations on recent LLMs demonstrate that LeanQuant is highly accurate, comparing favorably against recent competitive baselines in model quality, and scalable, achieving very accurate quantization of Llama-3.1 405B, one of the largest open-source LLMs to date, using two Quadro RTX 8000-48GB GPUs in 21 hours.
1 Introduction
Large language models (LLMs) have demonstrated impressive reasoning (Wei et al., 2022) and problem solving abilities (Kojima et al., 2022), and have shown the potential to bring transformative changes to various fields such as law (Kaddour et al., 2023), education (Kasneci et al., 2023), and medicine (Thirunavukarasu et al., 2023). However, deploying LLMs in a cost-effective manner presents significant challenges due to their substantial memory and computational demands (Chen et al., 2023), which hinders the accessibility and democratization of artificial intelligence (AI) (Kaddour et al., 2023).
Post-training quantization (PTQ) (Krishnamoorthi, 2018) is a technique for reducing the memory requirement for model inference by reducing the precision of floating-point weights of a pre-trained model and storing them in a compact low-bit-width format. PTQ offers the additional benefit of reducing the decoding latency of LLMs by reducing memory reads, since LLM inference is often bottlenecked by memory bandwidth (Kim et al., 2023). Although quantization causes a certain amount of precision loss in the parameters, the model quality can be reasonably preserved even in lower bit widths (Frantar et al., 2022; Chee et al., 2024). For many tasks, a quantized model is preferred over a full model due to its better size-accuracy trade-off (Dettmers & Zettlemoyer, 2023). As open-source foundational models continue to scale up in size (Dubey et al., 2024), accurate and efficient quantization becomes essential for making AI accessible to a wider audience. For instance, serving Llama-3.1 405B (Dubey et al., 2024) using its original 16-bit weights requires a cluster of 2 nodes, each with 8$\times$80GB GPUs, while the 4-bit quantized version can be deployed on a single nodes of 8$\times$48GB GPUs, eliminating the overhead of inter-node communication.
Challenges of Deploying Quantized Models One of the biggest challenges of successful deployment of quantized models is implementing optimized kernels for quantized GEMM (general matrix multiply) that are tailored to various hardware platforms and software frameworks. In order to accelerate inference of quantized models, fused kernels, which fuse dequantization and matrix multiplication in the same subroutine, have to be implemented and tuned for the specific hardware accelerator. These kernels require specialized designs and tunings for different hardware accelerators to be fully optimized (Park et al., 2022).
Recent quantization algorithms have chosen to employ specialized computations or custom data formats to reduce the impact of quantization on model quality, but they require more sophisticated kernel designs for efficient inference. For example, AQLM (Egiazarian et al., 2024) and QUIP# (Tseng et al., 2024) perform dequantization through look-ups from multi-dimensional or multi-bit codebooks, and Dotzel et al. (2024) proposed new data types such as Student Float to reduce quantization errors. While these approaches demonstrate promising results, their reliance on specialized operations and data formats can hinder their widespread adoption due to the need for optimized inference kernels for each hardware platform and software framework. For example, llama.cpp (Gerganov, 2023), a popular LLM inference engine that supports mobile devices, only supports affine and non-uniform quantization formats. Consequently, instead of focusing on developing better quantization methods with specialized operations, it may be more worthwhile to investigate improving the accuracy of existing widely adopted quantization formats, such as affine integer quantization and non-uniform quantization, which are supported by popular deep learning libraries (Paszke et al., 2019) and deployment frameworks (Kwon et al., 2023).
Scalability Challenges of Accurate Quantization To improve the quality of quantized models, existing approaches often incur higher computational overhead and require more hardware resources. As foundational models scale up in size (Hoffmann et al., 2022), these quantization approaches may struggle to scale to very large models such as Llama-3.1 405B (405 billion parameters) (Dubey et al., 2024). For instance, LLM-QAT (Liu et al., 2023) uses 100K samples of training data and hundreds of GPU-hours to recover the performance of a quantized LLaMA-13B model (Touvron et al., 2023a). For AQLM (Egiazarian et al., 2024), the time needed for quantizing a 7B to 70B LLM ranges from 1 to 14 days of an A100-80GB GPU. For SqueezeLLM (Kim et al., 2023), due to its use of the gradients of model parameters, quantizing a 70B LLM requires at least 240GB of total GPU memory, or 8$\times$32GB GPUs. As these accurate quantization approaches demand significant hardware resources and long optimization time, it is crucial to develop accurate methods that are efficient in terms of resource usage and time cost, to ensure the accessibility of increasingly large foundational models.
Our Proposal In this work, we propose LeanQuant, an accurate, versatile, and scalable quantization approach. We build upon the iterative loss-error-based quantization framework (Frantar & Alistarh, 2022; Frantar et al., 2022) and identify one of the biggest limitations of such methods: the min-max affine quantization grid introduces high loss errors due to the existence of outliers in the inverse Hessian diagonals. We introduce techniques for learning loss-error-aware quantization grids, which mitigate this issue and greatly improve the accuracy and quality of quantized models. We empirically demonstrate that LeanQuant compares favorably against competitive baselines in the 4/3/2-bit regions. Our approach is versatile, able to generalize to multiple commonly used quantization formats, such as affine and non-uniform quantization, allowing our quantized models to be directly compatible with existing highly optimized inference kernels (Frantar et al., 2024; Park et al., 2022) for maximum accessibility. Furthermore, our method is scalable and efficient. By designing and implementing a fused GPU kernel for LeanQuant grid learning, we achieve the accurate quantization of LLMs up to 123B in size using a single L40s-48GB GPU in 4 hours, and Llama-3.1 405B using 2 Quadro RTX 8000-48GB GPUs in 21 hours.
2 Background
In this section, we introduce the relevant background for our proposal including quantization grids and iterative loss-error-based quantization.
2.1 Quantization Grid
Quantization achieves model compression by representing the full-precision floating-point parameters with a limited set of grid points on the quantization grid. The number of points on the grid depends on the bit width: a $b$-bit code represents $2^{b}$ distinct values, so 2-bit quantization yields 4 grid points. The placement of these points significantly affects the precision of the original parameters, and imprecise placement can degrade model quality. To address this, different types of quantization grids, such as affine, non-uniform, and normal, have been introduced, which we survey as follows.
Affine Grid In an affine quantization grid (Krishnamoorthi, 2018), the grid points are evenly spaced over the dynamic range of a group of weights. The weights of the network are divided into groups of fixed size, e.g. every 128 contiguous parameters as a group. For min-max asymmetric affine quantization, the scaling factor $S$ and the zero-point $Z$ are recorded for each group of weights (the zero-point is omitted in the symmetric affine case). Then, the $i$-th weight $w_{i}$ in a group $\mathbf{w}$ is quantized to $b$-bit as follows,
$w_{i}^{\mathbf{int}}=\mathrm{clip}(\lfloor\frac{w_{i}}{S}\rceil+Z,0,2^{b}-1),\textrm{where }S=\frac{\max(\mathbf{w})-\min(\mathbf{w})}{2^{b}-1}\textrm{ and}Z=-\lfloor\frac{\min(\mathbf{w})}{S}\rceil$
$\mathrm{quant}_{\textit{aff}}(w_{i},S,Z)=(w_{i}^{\mathbf{int}}-Z)S$
[…]
Non-uniform Grid The grid points on a non-uniform grid are placed in a non-equidistant manner (Li et al., 2019). The motivation behind non-uniform quantization is to allow for finer precision in regions where model parameters are more concentrated or sensitive. Each row in a weight matrix has a distinct set of grid points $\mathcal{G}$, where $\lvert\mathcal{G}\rvert=2^{b}$ for $b$-bit non-uniform quantization. The weight $w_{i}$ is quantized to the nearest grid point in $\mathcal{G}$ as follows,
$\mathrm{quant}_{\textit{nu}}(w_{i},\mathcal{G})=\operatorname*{arg\,min}_{g\in\mathcal{G}}\lvert g-w_{i}\rvert$
Other Grid Types Previous works have observed that LLM parameters tend to be distributed similarly as Normal or Student T’s Distribution, hence they propose new grid types, such as NormalFloat (Dettmers et al., 2024) and Student Float (Dotzel et al., 2024), which place grid lines at the quantiles of these distributions. Our proposed method can also be extended to work with these quantization formats.
2.2 Iterative Loss-error-based Quantization
Iterative loss-error-based quantization (Frantar & Alistarh, 2022) is a promising framework for quantizing deep neural networks to low bit widths while maintaining strong performance on downstream tasks. In particular, Optimal Brain Quantization (OBQ) (Frantar & Alistarh, 2022), which is based on the seminal works by LeCun et al. (1989) and Hassibi et al. (1993), aims to minimize the impact of weight perturbations introduced by parameter quantization on the network’s task loss.
Let $\mathcal{L}(\mathbf{w}_{\mathcal{N}})$ be the task loss of a network $\mathcal{N}$ evaluated at the weights $\mathbf{w}_{\mathcal{N}}$ (flattened to a vector). Then, the OBQ objective is to minimize the loss error $\epsilon$, which is defined as
$\begin{split}\epsilon&=\mathcal{L}(\mathbf{w}_{\mathcal{N}}+\bm{\delta}_{\mathcal{N}})-\mathcal{L}(\mathbf{w}_{\mathcal{N}})\end{split}$
where $\bm{\delta}_{\mathcal{N}}$ is the weight perturbation introduced by quantization. The loss error $\epsilon$ can be approximated with a Taylor series (LeCun et al., 1989) as
$\epsilon=\underbrace{\big{(}\frac{\partial{\mathcal{L}}}{\partial{\mathbf{w}_{\mathcal{N}}}}\big{)}^{\top}\bm{\delta}_{\mathcal{N}}}_{\text{negligible}}+\frac{1}{2}\bm{\delta}_{\mathcal{N}}^{\top}\frac{\partial^{2}{\mathcal{L}}}{\partial{\mathbf{w}^{2}_{\mathcal{N}}}}\bm{\delta}_{\mathcal{N}}+\underbrace{O\big{(}\lVert\bm{\delta}_{\mathcal{N}}\rVert^{3}\big{)}}_{\text{negligible}}$
where the first term is omitted due to $\frac{\partial{\mathcal{L}}}{\partial{\mathbf{w}_{\mathcal{N}}}}\approx\mathbf{0}$ in a converged network, and the third and higher terms can be ignored due to small norms. Computing the exact Hessian $\mathbf{H}=\frac{\partial^{2}{\mathcal{L}}}{\partial{\mathbf{w}^{2}_{\mathcal{N}}}}$ in a deep network is difficult, hence OBQ leverages an approximation of loss error proposed by Nagel et al. (2020),
$\mathbb{E}(\epsilon)\approx\sum_{\mathbf{W}\in\mathcal{N}}\big{\lVert}\mathbf{W}\mathbf{X}-\hat{\mathbf{W}}\mathbf{X}\big{\rVert}_{F}^{2}$
where $\mathbf{W},\hat{\mathbf{W}},\mathbf{X}$ are the weight matrix, quantized weight matrix, and the input matrix to a linear layer in the network $\mathcal{N}$. As a result, the OBQ objective can be decomposed into layer-wise independent convex problems,
$\operatorname*{arg\,min}_{\hat{\mathbf{W}}}\lVert\mathbf{W}\mathbf{X}-\hat{\mathbf{W}}\mathbf{X}\rVert_{F}^{2}$ (1)
which can be further decomposed into row-wise independent problems, since Equation 1 can be written as a sum of squares over the rows of $\mathbf{W}$.
OBQ employs an iterative quantization approach, in which a single weight in a row $\mathbf{w}$ is quantized in each step, and then the remaining not-yet-quantized weights in the same row are updated to compensate for the introduced error. Given the constraint that the parameter $w_{i}$, indexed by $i$ in row $\mathbf{w}$, is being quantized, the optimal weight perturbation $\bm{\delta}$ to the remaining weights can be solved with the following Lagrangian,
$L(\bm{\delta},\lambda)=\frac{1}{2}\bm{\delta}^{\top}\mathbf{H}\bm{\delta}+\lambda\Big{(}\mathbf{e}_{i}^{\top}\bm{\delta}-\big{(}\mathrm{quant}(w_{i})-w_{i}\big{)}\Big{)}$ (2)
where the Hessian $\mathbf{H}=2\mathbf{X}\mathbf{X}^{\top}$ (from Equation 1) is computed on a small sample of input data and $e_{i}$ is the $i$-th standard basis vector. Solving Equation 2 yields the optimal weight perturbation $\bm{\delta}_{i}$ and the resulting loss error $\epsilon_{i}$ after quantizing $w_{i}$,
$\bm{\delta}_{i}=\frac{\mathrm{quant}(w_{i})-w_{i}}{\mathbf{H}^{-1}_{i,i}}\mathbf{H}^{-1}_{:,i},\;\;\epsilon_{i}=\frac{1}{2}\frac{\big{(}\mathrm{quant}(w_{i})-w_{i}\big{)}^{2}}{\mathbf{H}^{-1}_{i,i}}$ (3)
where $\mathbf{H}^{-1}_{i,i}$ and $\mathbf{H}^{-1}_{:,i}$ denotes the $i$-th diagonal entry and the $i$-th column of the inverse Hessian, respectively.
The loss error $\epsilon_{i}$ reflects the negative impact of quantizing parameter $w_{i}$ on the model quality, and it is always a non-negative value. OBQ utilizes the loss error $\epsilon_{i}$ as a heuristic for greedy optimization. Specifically, in each iteration, OBQ computes $\epsilon$ for all weights in a row and greedily selects the $i$-th parameter with the minimum $\epsilon_{i}$ to quantize. Then, $w_{i}$ is round to the nearest value on the quantization grid, and the remaining weights are updated via $\mathbf{w}\leftarrow\mathbf{w}-\bm{\delta}_{i}$, and the updated inverse Hessian for the remaining weights, with the $i$-th row and column removed from $\mathbf{H}$, is computed as
${\mathbf{H}^{-1}_{-i,-i}=\big{(}\mathbf{H}^{-1}-\frac{\mathbf{H}^{-1}_{:,i}\mathbf{H}^{-1}_{i,:}}{\mathbf{H}^{-1}_{i,i}}\big{)}_{-i,-i}}$ (4)
This iterative process continues until all weights are quantized.
Scaling to Billion-Parameter LLMs Using Cholesky and Dampening OBQ produces accurate post-training quantized models for million-parameter networks, but fails to scale to billion-parameter LLMs due to two primary reasons: the inefficient time complexity and the accumulation of numerical inaccuracies during updates. To improve its computational efficiency, Frantar et al. (2022) propose to quantize the weights in a fixed non-greedy order for all rows, and keep the weight updates within a block of $B$ columns at a time. To prevent model quality collapse resulted from the accumulation of numerical inaccuracies by repeated weight updates, Frantar et al. (2022) propose to apply a mild dampening (1% of the average diagonals) to the diagonal entries of the Hessian $\mathbf{H}$ and leverage a Cholesky decomposition of the inverse Hessian $\mathbf{H}^{-1}$ in place of the update in Equation 4. The resulting algorithm is GPTQ, which is able to efficiently quantizes billion-parameter LLMs.
3 Methodology
Figure 1: (Left) The empirical distributions of inverse Hessian diagonals, computed on 262K tokens from the C4 dataset for the Llama-3-8B model, contain outliers that can cause high loss errors. (Right) Our proposed loss-error-aware non-uniform and affine grids better preserves the quantized precision of outliers, leading to more accurate quantized models.
In this section, we introduce our proposed approach Loss-error-aware network Quantization (LeanQuant), for accurately and efficiently quantizing LLMs.
3.1 Revisiting the Loss Error
To motivate our proposed approach, we first revisit the loss error $\epsilon_{i}$ in Equation 3, which approximates the (detrimental) increase in the network’s task loss, introduced by quantizing weight $w_{i}$. This error $\epsilon_{i}$ has been used as a heuristic in multiple previous works (LeCun et al., 1989; Hassibi et al., 1993; Singh & Alistarh, 2020; Frantar & Alistarh, 2022) for choosing the next best weight $i$ to prune or quantize. It has been shown to be a highly informative metric for measuring the impact of quantization.
By examining Equation 3, one finds that the loss error $\epsilon_{i}$ is proportional to the square of weight quantization error and inversely proportional to the diagonal entry of the inverse Hessian, i.e.,
$\epsilon_{i}\propto\big{(}\mathrm{quant}(w_{i})-w_{i}\big{)}^{2}\;\text{ and }\;\epsilon_{i}\propto\frac{1}{\mathbf{H}^{-1}_{i,i}}$ (5)
Hence, we further examine the empirical distribution of $\frac{1}{\mathrm{diag}(\mathbf{H}^{-1})}$, to which the loss error row vector $\bm{\epsilon}$ is proportional. We obtain the empirical distributions on layers of Llama-3-8B (Dubey et al., 2024) with 128 sequences of length 2048 tokens from the C4 dataset (Raffel et al., 2020), and compute the inverse Hessian as $\mathbf{H}^{-1}=(2\mathbf{X}\mathbf{X}^{\top})^{-1}$ where $\mathbf{X}$ is the layer input matrix. As shown in Figure 1, The majority of the inverse diagonals are concentrated in low-magnitude regions, with a few outliers having high magnitudes. Quantizing the weights corresponding to these outliers can lead to high loss errors if these weights are not well-aligned with the quantization grid points. Preserving the quantized precision of the weights corresponding to these inverse-diagonal outliers is especially important because the loss error increases quadratically with their quantization error (Equation 5). Iterative loss-error-based quantization approaches (OBQ, GPTQ, etc.) employ min-max affine quantization grid, which is suboptimal for preserving the quantized precision of the inverse-diagonal outliers, leading to high loss errors and model quality degradation. Our idea is to learn quantization grids that minimize the loss error $\epsilon$.
3.2 Loss-Error-Aware Network Quantization
The min-max affine grid employed by previous iterative loss-error-based quantization methods do not account for the existence of inverse-diagonal outliers, which may cause significant model quality degradation. We propose loss-error-aware quantization grids, which preserve the quantized precision of weights corresponding to the inverse diagonal outliers, to produce more accurate models. We propose techniques to learn loss-error-aware grids for different quantization formats, including non-uniform and affine, and design fused GPU kernels to achieve efficient and scalable quantization.
3.2.1 Non-Uniform Loss-Error-Aware Grid
For non-uniform quantization, we perform clustering with the model parameters, weighted by their corresponding exponentiated inverse Hessian diagonals, to obtain a set of grid points that are loss-error-aware. Concretely, we learn the set of grid points $\mathcal{G}$ for $b$-bit quantization through the following objective,
$\operatorname*{arg\,min}_{\mathcal{G}:\left|\mathcal{G}\right|=2^{b}}\sum_{i}(\mathbf{H}^{-1}_{i,i})^{-p}\left|\mathrm{quant}_{\textit{nu}}(w_{i},\mathcal{G})-w_{i}\right|^{2}$ (6)
where $p$ is a hyperparameter that indicates the strength of preserving the quantized precision of outliers. We leverage k-means (Lloyd, 1982) to optimize the objective. After the quantization grid $\mathcal{G}$ has been established, the weights are iteratively quantized to the nearest grid point in $\mathcal{G}$.
Grid Initialization The quality of clustering results is highly dependent on initialization (Arthur et al., 2007), since Lloyd’s Algorithm (Lloyd, 1982) converges to a locally optimal solution. Hence in lower bit width regions, the initialization of grid points for clustering can greatly influence the quantized model quality. K-means++ (Arthur et al., 2007) leverages iterative sampling to select initial centers that are likely far from each other to obtain more optimal results. However, we found that it still tends to produce sub-optimal results in lower bit width regions (3 and 2-bit). We propose a lightweight and robust grid point initialization method: uniformly spaced grid initialization, which uses uniformly spaced points $\mathcal{G}_{\textrm{init}}$ over the dynamic range of the weights as initial centers, where
$\mathcal{G}_{\textrm{init}}=\Big{\{}\min(\mathbf{w})+\frac{\max(\mathbf{w})-\min(\mathbf{w})}{2^{b}-1}t\,\Big{|}\,t\in\{0,\dots,2^{b}-1\}\Big{\}}$
We perform ablative experiments to study the effectiveness of uniformly spaced grid initialization in Table 11 in the Appendix.
3.2.2 Loss-error-aware Affine Grid
We propose a different strategy for learning affine loss-error-aware grid than non-uniform grid. The challenge with learning affine grid is that the grid points are uniformly placed over an interval, so the clustering strategy does not apply. Searching for the optimal scaling factor $S$ and zero-point $Z$ over the real numbers can be done using gradient descent, but it is potentially time and memory consuming and prone to finding local minimas. Hence, we adopt the strategy of searching for the optimal $S$ and $Z$ in a constrained space $\mathbb{S}$, and formulate the problem of learning affine loss-error-aware grid as
[…]
where $R=\max(\mathbf{w})-\min(\mathbf{w})$ is the dynamic range of the weights, and $T$ is the number of partitions over $R$. To obtain the $S$ and $Z$ that minimize the loss errors, we enumerate candidates for $S$ and $Z$ by shrinking the min-max range of the weights in discrete steps. The value of $T$ determines the granularity of the search, which we set to 2048.
Efficient Fused GPU Kernel for Grid Learning The learning process in Equation 7 is costly due to the enumeration of $(\frac{T}{2})^{2}$ pairs of scaling factor and zero-point. Simply looping over all enumerations is time-consuming. Hence, we design and implement a fused kernel that leverages the parallel processing power of GPUs for efficient grid searching. We designate each thread block to be responsible for processing a single group of weights, and each thread in the thread block to be responsible for enumerating all combinations of a single $t_{\mathrm{min}}$ and all $t_{\mathrm{max}}$. The minimum loss for each $t_{\mathrm{min}}$ is accumulated by each thread, and they are aggregated at the end on the thread block level to determine the optimal scaling factor and zero-point for the current weight group. This mode of processing enables the parallel computation of the optimal scaling factor and zero-point for all weight groups, accelerating the end-to-end quantization process by more than 50$\times$. An efficiency analysis of our fused kernel is presented in Section 4.3.
3.2.3 LeanQuant
Our proposed loss-error-aware quantization grid can integrate with any iterative loss-error-based quantization approaches to improve quantized model quality. A comparison between min-max affine quantization grid with loss-error-aware non-uniform and affine grid learned for a layer of Llama-3-8B (Dubey et al., 2024) is presented in Figure 1. We propose LeanQuant, which integrates loss-error-aware grid with GPTQ (Frantar et al., 2022), and present it in Algorithm 1. Additionally, for quantizing million-parameter models more accurately, we propose LeanQuant-Exact, which integrates loss-error-aware grid with OBQ (Frantar & Alistarh, 2022), and present it in Algorithm 2 in the Appendix. We further denote the underlying grid type used by LeanQuant with subscripts, such as LeanQuantaff and LeanQuantnu for affine and non-uniform grid, respectively.
The parameter $p$ balances the strength of precision preservation between inverse-diagonal outliers and non-outliers. Higher values of $p$ prioritizes the precision preservation of outliers more, while $p=0$ prioritizes the preservation of all weights equally. We found the quantized model quality to be not very sensitive to values of $p$, and use $p=4$ for all models. A sensitivity analysis for $p$ is given in Section 4.3.
Algorithm 1 LeanQuant for LLM quantization
1:Input: weight matrix $\mathbf{W}\in\mathbb{R}^{r\times c}$, input matrix $\mathbf{X}$, bit width $b$, block size $B$, dampening factor $df$, outlier preservation strength $p$
2:Output: Quantized matrix $\hat{\mathbf{W}}$
3:$\hat{\mathbf{W}}\leftarrow\mathbf{0}_{r\times c}$
4:$\mathbf{E}\leftarrow\mathbf{0}_{r\times B}$
5:$\mathbf{H}\leftarrow 2\mathbf{X}\mathbf{X}^{\top}$
6:$\mathbf{H}^{-1}\leftarrow\mathrm{Cholesky}\Big{(}\big{[}\mathbf{H}+df\cdot\mathrm{avg}\big{(}\mathrm{diag}(\mathbf{H})\big{)}\cdot\mathbf{I}\big{]}^{-1}\Big{)}$triangleright apply dampening, inversion, and Cholesky decomposition
7:if using non-uniform grid then
8:   $\mathcal{G}_{k}\leftarrow\operatorname*{arg\,min}\limits_{\mathcal{G}:\left|\mathcal{G}\right|=2^{b}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{nu}}(\mathbf{W}_{k,:},\mathcal{G})-\mathbf{W}_{k,:}\big{|}^{2}$ forall $k\in\{0,\dots,r-1\}$triangleright E.6
9:else if using affine grid then
10:   $S_{k},Z_{k}\leftarrow\operatorname*{arg\,min}\limits_{(S,Z)\in\mathbb{S}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{aff}}(\mathbf{W}_{k,:},\mathcal{S},Z)-\mathbf{W}_{k,:}\big{|}^{2}$ forall $k\in\{0,\dots,r-1\}$triangleright E.7
11:end if
12:for $i\leftarrow 0,B,2B,\dots$ do $\triangleright$ apply block-wise quantization
13:   for $j\leftarrow i,\dots,i+B-1$ do
14:      if using non-uniform grid then
15:         $\hat{\mathbf{W}}_{k,j}\leftarrow\mathrm{quant}_{\textit{nu}}(\mathbf{W}_{k,j},\mathcal{G}_{k})$ forall $k\in\{0,\dots,r-1\}$triangleright quantize to non-uniform grid
16:      else if using affine grid then
17:         $\hat{\mathbf{W}}_{k,j}\leftarrow\mathrm{quant}_{\textit{aff}}(\mathbf{W}_{k,j},S_{k},Z_{k})$ forall $k\in\{0,\dots,r-1\}$triangleright quantize to affine grid
18:      end if
19:      $\mathbf{E}_{:,j-1}\leftarrow\frac{\mathbf{W}_{:,j}-\hat{\mathbf{W}}_{:,j}}{\mathbf{H}^{-1}_{j,j}}$
20:      $\mathbf{W}_{:,j:(i+B)}\leftarrow\mathbf{W}_{:,j:(i+B)}-\mathbf{E}_{:,j-i}\cdot\mathbf{H}^{-1}_{j,j:(i+B)}$
21:   end for
22:   $\mathbf{W}_{:,(i+B):}\leftarrow\mathbf{W}_{:,(i+B):}-\mathbf{E}\cdot\mathbf{H}^{-1}_{i:(i+B),(i+B):}$
23:end for
24:return $\hat{\mathbf{W}}$
4 Experiments
Table 1: Zero-shot accuracy of quantized LLMs on benchmarks. The results of more models can be found in Table 8 of the Appendix. 2-bit quantization is unsupported by the SqueezeLLM codebase.
Method Bits ARC LAMBADA MMLU HellaS PIQA WinoG Avg.
Easy Chg Std OpenAI STEM Human. Social Other
Llama-3-8B
BF16 16 80.30 50.17 68.85 75.82 53.82 54.88 73.29 70.42 60.11 79.71 73.56 67.36
GPTQ 4.00 74.83 44.11 63.42 70.75 47.29 52.28 66.04 64.89 57.98 77.26 71.82 61.58
Affine OmniQuant 4.00 76.89 47.35 61.05 69.16 49.38 49.05 66.62 64.40 58.25 78.84 71.98 63.00
LeanQuantaff 4.00 76.60 46.93 66.89 74.07 51.89 52.96 70.04 68.43 58.47 77.91 72.77 65.18
GPTQ 3.00 50.84 24.32 24.16 38.89 26.23 29.16 34.38 30.00 45.07 64.64 60.69 37.75
OmniQuant 3.00 60.90 30.12 21.08 27.63 26.32 27.80 29.51 29.90 46.98 68.17 59.98 38.95
LeanQuantaff 3.00 69.44 35.75 46.81 65.42 42.59 44.78 58.17 56.97 52.72 74.86 69.93 56.13
GPTQ 2.00 25.46 22.53 0.00 0.00 21.06 23.95 21.16 23.78 25.66 52.77 51.54 24.25
OmniQuant 2.00 26.81 21.67 0.00 0.00 21.34 24.21 21.71 23.98 25.90 53.75 47.43 24.26
LeanQuantaff 2.00 35.06 18.26 11.33 14.71 21.31 24.17 21.71 24.01 31.43 59.30 51.85 28.47
Non-uniform SqueezeLLM 4.05 79.59 49.32 66.18 73.24 51.13 53.32 70.78 68.59 59.10 79.33 73.80 65.85
LeanQuantnu 4.05 79.50 49.15 67.36 74.95 52.17 53.16 71.40 68.75 59.19 78.89 74.11 66.24
SqueezeLLM 3.02 73.19 43.52 58.22 66.58 43.61 46.57 61.91 60.03 56.17 77.64 69.22 59.70
LeanQuantnu 3.02 77.74 47.01 63.32 72.17 48.84 49.05 65.45 62.79 56.42 78.24 71.67 62.97
SqueezeLLM 2.01 - N/A -
LeanQuantnu 2.01 58.21 26.62 31.22 39.16 25.98 25.48 27.01 26.65 40.78 68.01 60.38 39.05
Llama-2-7B
FP16 16 76.26 43.43 68.33 73.88 34.38 39.79 47.32 47.12 57.10 78.07 68.98 57.70
GPTQ 4.00 74.16 40.78 65.38 71.94 32.67 36.92 42.61 42.61 55.99 77.48 68.32 53.47
Affine OmniQuant 4.00 74.12 40.70 64.10 70.62 28.80 32.18 34.71 35.79 55.37 76.93 68.67 52.91
LeanQuantaff 4.00 75.00 41.21 65.03 72.02 34.82 36.94 46.77 44.54 55.32 77.15 68.75 56.14
GPTQ 3.00 66.29 34.22 46.46 58.18 28.20 26.99 32.11 29.90 49.05 73.23 62.83 44.12
OmniQuant 3.00 70.12 37.29 53.27 66.66 29.05 31.05 30.61 30.38 52.58 74.05 66.46 49.23
LeanQuantaff 3.00 71.84 38.99 59.13 69.05 33.56 32.96 41.11 40.10 52.19 75.57 66.69 52.84
GPTQ 2.00 25.97 21.67 0.00 0.00 21.31 23.25 21.11 23.01 25.76 51.74 48.78 23.66
OmniQuant 2.00 37.42 21.76 1.28 3.24 21.47 24.14 21.74 23.91 29.59 57.18 51.93 26.70
LeanQuantaff 2.00 41.08 20.99 16.98 21.93 21.25 24.06 21.77 23.88 31.94 61.64 56.51 31.09
Non-uniform SqueezeLLM 4.05 75.59 41.98 67.81 72.79 34.32 38.94 45.40 44.96 56.80 77.48 68.43 56.77
LeanQuantnu 4.05 75.97 42.66 68.14 74.25 34.35 39.06 46.05 46.51 56.03 77.86 69.38 57.30
SqueezeLLM 3.02 73.06 40.27 61.96 70.11 33.75 35.22 43.35 43.16 54.15 76.50 67.88 54.49
LeanQuantnu 3.02 73.74 40.19 66.12 73.16 32.25 35.54 43.40 43.39 53.24 76.44 68.35 55.07
SqueezeLLM 2.01 - N/A -
LeanQuantnu 2.01 51.81 23.98 28.68 38.21 22.26 23.89 22.49 24.01 35.88 66.38 58.17 35.98
Mistral-7B
BF16 16 80.77 50.09 69.38 75.63 50.46 53.48 69.35 68.01 61.26 80.58 73.88 66.62
GPTQ 4.00 79.00 46.25 66.99 73.67 46.24 50.82 66.20 64.66 59.36 79.65 72.93 62.68
Affine OmniQuant 4.00 78.49 46.25 63.28 71.20 45.96 51.35 65.68 64.76 60.19 79.87 71.90 63.54
LeanQuantaff 4.00 79.71 48.04 68.33 75.70 47.42 51.84 68.05 66.43 59.65 80.41 73.48 65.37
GPTQ 3.00 70.54 38.65 52.63 62.10 36.31 38.89 49.20 47.86 54.76 77.58 67.96 52.60
OmniQuant 3.00 70.54 35.07 35.49 46.54 33.71 32.88 40.23 37.85 52.35 75.19 63.93 47.62
LeanQuantaff 3.00 76.94 44.62 65.63 74.60 44.18 45.59 61.20 59.03 56.36 78.89 72.30 61.76
GPTQ 2.00 26.73 22.27 0.00 0.00 23.31 24.46 23.86 23.42 25.35 51.52 49.72 24.39
OmniQuant 2.00 27.06 21.67 0.00 0.00 21.25 24.29 21.71 23.98 25.89 51.25 51.54 24.42
LeanQuantaff 2.00 57.91 27.22 36.91 49.00 24.23 24.91 24.60 27.29 40.27 69.15 60.46 40.18
Non-uniform SqueezeLLM 4.05 79.73 49.06 68.28 74.93 48.81 52.73 68.87 66.98 59.80 80.25 73.56 65.73
LeanQuantnu 4.05 79.80 48.89 69.03 76.03 48.84 52.86 68.87 66.69 60.19 80.14 74.59 65.99
SqueezeLLM 3.02 77.54 45.93 64.06 71.43 43.96 47.93 62.69 59.16 58.76 79.43 71.98 62.08
LeanQuantnu 3.02 77.74 45.99 67.59 76.07 44.24 47.97 62.14 62.47 57.28 79.27 72.22 63.00
SqueezeLLM 2.01 - N/A -
LeanQuantnu 2.01 63.47 30.55 41.01 54.61 31.34 29.97 32.14 33.96 42.29 71.38 64.01 44.97
In this section, we perform extensive experiments to validate the effectiveness and scalability of our proposed LeanQuant for quantizing LLMs against competitive baselines. We first introduce the baselines, models, evaluation metrics and datasets, and hardware used for the experiments. We then describe the experimental results and findings, and analyze the efficiency and scalability of our proposal. Finally, we perform ablative experiments to validate each component of proposed approach.
Baselines We compare LeanQuantaff against competitive affine quantization approaches AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), and OmniQuant (Shao et al., 2024), and LeanQuantnu against the existing state-of-the-art non-uniform method SqueezeLLM (Kim et al., 2023). For the baselines, we use the quantized models provided by their official repository where possible, and quantize the unavailable models using their official codebase and recommended hyperparameters. More details on baseline reproduction and evaluation methods can be found in Section C of the Appendix. For all LeanQuant models, we use a small calibration set of 128 sequences of 2048 tokens from the C4 dataset (Raffel et al., 2020) for computing the Hessian $\mathbf{H}$, and set $p=4$.
Models We consider the following recent, popular LLMs for quantization: Llama 1/2/3 series models (Touvron et al., 2023a; b; Dubey et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023), Mistral-Large-Instruct-2407 (123B) (Mistral AI Team, 2024), and Llama-3.1-405B-Instruct (Dubey et al., 2024).
Evaluation Metrics and Datasets We evaluate quantized LLMs using the perplexity metric on the datasets WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), and zero-shot accuracy on the benchmarks ARC (Clark et al., 2018), LAMBADA (Paperno et al., 2016), MMLU (Hendrycks et al., 2020), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), and WinoGrande (Sakaguchi et al., 2021). We also quantize and evaluate the instruction-following Llama-3-8B-Instruct using OpenAI GPT-4o (2024-05-13) as a judge on the MT-Bench (Zheng et al., 2023), and the results are presented in Section E in the Appendix.
Testbed Hardware LeanQuant models are quantized using a machine quipped with an L40s-48GB GPU, an AMD EPYC 7R13 48-Core CPU, and 370GB of RAM. To fit Llama-3.1-405B-Instruct in RAM, which is around 800GB in size, we use a machine equipped with 2 Quadro RTX 8000 GPUs, an AMD EPYC 7742 64-Core CPU, and 1.48TB of RAM.
Table 2: Zero-shot accuracy comparison of the quantized 123B Mistral-Large-Instruct-2407 model.
Model Method Bits Arc LAMBADA MMLU Avg.
Easy Chg. Std. OpenAI STEM Human. Social Other
Mistral-Large-Instruct-2407 GPTQ 4.00 84.60 63.99 74.38 80.52 76.31 77.23 89.31 85.23 78.95
LeanQuantaff 4.00 85.14 63.99 74.99 81.14 76.56 77.32 89.21 85.87 79.28
Table 3: Zero-shot accuracy comparison of the quantized Llama-3.1-405B-Instruct model.
Model Method Group Size Bits Arc-E Arc-C LAMBADA-Std PIQA Avg
Llama-3.1-405B-Instruct GPTQ 128 4.25 88.21 65.10 76.96 82.75 78.26
LeanQuantaff 128 4.25 88.34 65.70 77.86 83.03 78.73
4.1 Main Results
Accuracy and Perplexity The zero-shot accuracy of quantized models on benchmarks are presented in Table 1, as well as in Table 8 in the Appendix, and the perplexity results are shown in Table 7 in the Appendix. At the same bit width, LeanQuant achieves significantly better (lower) perplexity than GPTQ and AWQ, and performs on par with OmniQuant and SqueezeLLM. However, perplexity may not be a representative metric for evaluating the accuracy of quantized models. In terms of zero-shot accuracy on various benchmarks, LeanQuantaff mostly outperforms GPTQ and OmniQuant, and LeanQuantnu similarly performs better than SqueezeLLM in most cases. We highlight that LeanQuantaff improves the average zero-shot accuracy on 11 tasks over OmniQuant by 17.18% for 3-bit Llama-3-8B, and by 14.14% for 3-bit Mistral-7B. Compared to GPTQ, LeanQuantaff improves the average zero-shot accuracy by 18.38% for 3-bit Llama-3-8B, and by 9.16% for 3-bit Mistral-7B.
Effectiveness on Very Large LLMs We quantize the 123B Mistral-Large-Instruct-2407 and the 405B Llama-3.1 model using LeanQuantaff and GPTQ, and present their zero-shot accuracy in Table 2 and 3, respectively. OmniQuant and SqueezeLLM fail to quantize to these models due to GPU out-of-memory errors. For Llama-3.1 405B, we use a smaller set of evaluation tasks due to the high inference costs. LeanQuantaff models mostly outperforms GPTQ in zero-shot accuracy. We employ row-wise quantization for Mistral-Large and group-wise quantization (with size 128) for Llama-3.1. This showcases that our method is effective for both row-wise affine quantization and group-wise affine quantization.
4.2 Memory and Time Efficiency
We report the maximum GPU memory consumption of LeanQuant and the baselines during quantization on models of different sizes in Table 5. LeanQuant is significantly more memory efficient than OmniQuant and SqueezeLLM: it successfully scales to 123B Mistral-Large using a single 48GB GPU, and to 405B Llama-3.1 models using two 48GB GPUs, while OmniQuant fails to quantize Llama-3-70B and SqueezeLLM fails to quantize Llama-3-8B on a single 48GB GPU. The time cost of LeanQuant for different sized models are reported in Table 9 in the Appendix. LeanQuant can quantize 7B/8B models in less than an hour, the 123B model in 4.2 hours, and the 405B model in 20.7 hours.
4.3 Ablation Study
Q1: Does LeanQuant effectively reduce the loss error $\epsilon$ compared to other iterative loss-error-based methods? Yes, LeanQuant effectively reduces loss errors $\epsilon$ compared to GPTQ, as shown in Figure 2, as well as in Figure 4 in the Appendix. The sum of loss errors are computed as Equation 3. Moreover, non-uniform LeanQuant generally achieves lower loss errors than affine LeanQuant, due to more degrees of freedom in the grid point placements, which also explains why LeanQuantnu achieves higher accuracy than LeanQuantaff on benchmarks in Table 1.
Q2: Is LeanQuant sensitive to the hyperparameter $p$? No, we found LeanQuant to be not very sensitive to $p$. A sensitivity analysis on the hyperparameter $p$ is given in Table 10 in the Appendix. LeanQuant works well with $p$ values of 3 or 4.
Q3: Is uniformly spaced grid initialization beneficial for model quality? Yes, uniformly spaced grid initialization consistently outperforms k-means++ (Arthur et al., 2007) initialization on different models in 3-bit and 2-bit regions, as shown in Table 11 in the Appendix.
Q4: Does the fused GPU kernel for LeanQuantaff accelerate quantization? Yes, our fused kernel for learning affine grids accelerate the end-to-end quantization process by more than 50$\times$, as shown in Table 5, which enables LeanQuant to be scaled to very large models.
Figure 2: Comparison of loss errors $\epsilon$, summed over each layer, for GPTQ and LeanQuant (affine and non-uniform) during iterative quantization.
Table 4: Peak GPU memory consumption of different algorithms during 4-bit quantization. “OOM” indicates out of memory on a single 48GB GPU, except for Llama-3.1-405B where we use 2 48GB GPUs.
Model OmniQuant SqueezeLLM LeanQuant
Llama-3-8B 25.3 GB OOM 7.9 GB
Llama-3-70B OOM OOM 17.2 GB
Mistral-Large (123B) OOM OOM 33.0 GB
Llama-3.1-405B OOM OOM 65.4 GB
Table 5: Comparison of total time needed for quantizing Llama-3-8B with and without our fused kernel for loss-error-aware affine grid learning.
Fused Kernel Group Size Bits Quant. Time
- 4.00 15.1 hrs
- 4.00 0.27 hrs
128 4.25 >100 hrs
128 4.25 0.40 hrs
5 Related Works
Iterative Loss-error-based Compression Optimal Brain Damage (LeCun et al., 1989) introduced a saliency-score-based iterative pruning algorithm for neural networks, and Optimal Brain Surgeon (Hassibi & Stork, 1992; Hassibi et al., 1993) extended it to apply a weight update to compensate for the error introduced in each iteration. These methods inspired a number of works on model pruning (Guo et al., 2016; Singh & Alistarh, 2020; Yu et al., 2022) and weight quantization (Li et al., 2021; Frantar & Alistarh, 2022; Frantar et al., 2022).
Efficient LLM Inference LLM inference is computationally and memory demanding, and existing works accelerate inference and reduce memory requirements through post-training weight quantization (Dettmers et al., 2022; Lin et al., 2024; Frantar et al., 2022; Chee et al., 2024; Kim et al., 2023; Shao et al., 2024; Egiazarian et al., 2024; Tseng et al., 2024), pruning (Frantar & Alistarh, 2023; Ashkboos et al., 2024), weight-activation quantization (Xiao et al., 2023), offloading Sheng et al. (2023), etc.
6 Conclusion
In this work, we propose LeanQuant, an accurate, versatile, and scalable quantization method for LLMs. Motivated by the finding that the min-max affine grid causes large errors in the network’s task loss in iterative loss-error-based methods, we propose to learn loss-error-aware grids to enable more accurate quantized models, and design fused kernels for efficient and scalable quantization. Our method generalizes to multiple quantization formats to enable greater accessibility. Extensive empirical evaluations reveal that our quantized models compares favorably against competitive baselines in accuracy, and can scale to Llama-3.1 405B, one of the largest open-source LLM to date.
Appendix
Appendix A LeanQuant-Exact
The pseudocode of LeanQuant-Exact for accurately quantizing million-parameter networks is presented in Algorithm 2.
Algorithm 2 LeanQuant-Exact for Millon-parameter Networks
1:Input: a row $\mathbf{w}\in\mathbb{R}^{c}$ in the weight matrix, sample input matrix $\mathbf{X}$, bit width $b$, hyperparameter $p$
2:Output: Quantized row $\hat{\mathbf{w}}$
3:$\hat{\mathbf{w}}\leftarrow\mathbf{0}_{c}$
4:$\mathbf{H}^{-1}\leftarrow(2\mathbf{X}\mathbf{X}^{\top})^{-1}$
5:if using non-uniform grid then
6:   $\mathcal{G}\leftarrow\operatorname*{arg\,min}\limits_{\mathcal{G}:\left|\mathcal{G}\right|=2^{b}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{nu}}(\mathbf{w},\mathcal{G})-\mathbf{w}\big{|}^{2}$triangleright E. 6
7:else if using affine grid then
8:   $S,Z\leftarrow\operatorname*{arg\,min}\limits_{(S,Z)\in\mathbb{S}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{aff}}(\mathbf{w},\mathcal{S},Z)-\mathbf{w}\big{|}^{2}$triangleright E. 7
9:end if
10:for $j\leftarrow 1,\dots,c$ do
11:   if using non-uniform grid then
12:      $i\leftarrow\operatorname*{arg\,min}_{i}\frac{(\mathrm{quant}_{\textit{nu}}(w_{i},\mathcal{G})-w_{i})^{2}}{2\mathbf{H}^{-1}_{i,i}}$
13:      $\hat{w}_{i}\leftarrow\mathrm{quant}_{\textit{nu}}(w_{i},\mathcal{G})$
14:   else if using affine grid then
15:      $i\leftarrow\operatorname*{arg\,min}_{i}\frac{(\mathrm{quant}_{\textit{aff}}(w_{i},S,Z)-w_{i})^{2}}{2\mathbf{H}^{-1}_{i,i}}$
16:      $\hat{w}_{i}\leftarrow\mathrm{quant}_{\textit{aff}}(w_{i},S,Z)$
17:   end if
18:   $\mathbf{w}\leftarrow\mathbf{w}-\frac{\mathbf{H}^{-1}_{:,i}}{\mathbf{H}^{-1}_{i,i}}\big{(}w_{i}-\hat{w}_{i}\big{)}$
19:   $\mathbf{H}^{-1}\leftarrow\mathbf{H}^{-1}-\frac{\mathbf{H}^{-1}_{:,i}\mathbf{H}^{-1}_{i,:}}{\mathbf{H}^{-1}_{i,i}}$
20:end for
21:return $\hat{\mathbf{w}}$
A.1 BERT Experiments with LeanQuant-Exact
We compare the performance of BERT models (Devlin et al., 2018), quantized with OBQ (Frantar & Alistarh, 2022) and LeanQuantnu-Exact, on the SQuAD dataset (Rajpurkar et al., 2016). We quantize the 12-layer BERT-base (Devlin et al., 2018) and the 3-layer BERT-3 variant from Kurtic et al. (2022) to 3 and 4 bits. OBQ and LeanQuant-Exact are calibrated using 1024 samples from the training set, and the F1 score is reported on the test set.
Method Bits BERT-3 BERT
FP32 32 84.66 88.53
OBQ 4.03 84.40 87.96
LeanQuantnu-Exact 4.13 84.58 88.49
OBQ 3.03 83.47 84.72
LeanQuantnu-Exact 3.06 84.20 86.21
Table 6: F1 scores on SQuAD of BERT models quantized using OBQ and LeanQuantnu-Exact. LeanQuantnu-Exact outperforms OBQ in maintaining model quality.
Appendix B Discussion on Error Accumulation During Iterative Quantization
LeanQuant prevents drastic increase to the task loss by learning the quantization grid for better preservation of the precision of outlier inverse diagonals. However, since the not-yet-quantized weights will shift during the iterative quantization process and the quantization grid is fixed beforehand, one potential problem arises: the quantization grid is no longer well-aligned with the outliers after certain iterations. Fortunately, this is not a problem in practice. The loss-error-awareness property of LeanQuant grids prevents high-norm weight perturbations $\bm{\delta}_{i}$ (Equation 3) from ocurring, hence the weights do not shift by much during the iterations. Furthermore, no new inverse-diagonal outliers will arise during the iterative quantization process, since the remaining inverse diagonals only decrease in magnitude towards zero after each column and row removal (Equation 4).
Appendix C Experiment Details
Baseline Reproduction
We use the quantized models provided by the official repository where possible. We obtained quantized LLaMA-7B, LLaMA-13B, Llama-2-7B, Llama-2-13B from the OmniQuant repository, and LLaMA-7B, LLaMA-13B, Llama-2-7B, Llama-2-13B, Mistral-7B from the SqueezeLLM repository. We obtained the community-driven GPTQ-quantized version of Llama-3.1-405B-Instruct from HuggingFace  [^1] https://huggingface.co/hugging-quants/Meta-Llama-3.1-405B-Instruct-GPTQ-INT4. The other quantized models are reproduced using the official codebases and recommended hyperparameters. For OmniQuant, we set the training epochs to 20, enable Learnable Weight Clipping (LWC), set an LWC learning rate of 1e-2. For SqueezeLLM, there is no tunable parameters. For GPTQ, we turn on activation ordering (quantizing columns in order of decreasing activation size) for more accurate model.
Perplexity Evaluations
We follow the perplexity evaluation procedure described by (Frantar et al., 2022): sequences from the test set of the WikiText2 and C4 datasets (Merity et al., 2016; Raffel et al., 2020) are concatenated into 128 sequences of length 2048 tokens for perplexity testing.
Accuracy Evaluations
We use lm-evaluation-harness (Gao et al., 2023) for evaluating zero-shot accuracy on tasks. The task names we evaluate are lambada, ai2_arc, winogrande, piqa, hellaswag, mmlu.
Appendix D Perplexity Evaluations
The perplexity evaluation results on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) for quantized models are presented in Table 7.
Appendix E LLM-as-a-Judge
LLM as a Judge The evaluation results on MT-Bench using GPT-4o (2024-05-13) as a judge are presented in Figure 3. We pitch 3-bit and 4-bit, with group size of 128, LeanQuantaff against OmniQuant, and 4-bit LeanQuantnu against SqueezeLLM. LeanQuant achieves higher win rate than the baselines.
Figure 3: Evaluation of quantized Llama-3-8B-Instruct on MT-Bench using OpenAI GPT-4o as a judge. The win rates reported exclude ties.
Appendix F Accuracy Results on More Models
The zero-shot accuracy results on benchmarks for quantized LLaMA-7B, LLaMA-13B, Llama-2-7B (Touvron et al., 2023a; b) are presented in Table 8.
Table 7: Perplexity evaluations of Llama models under different quantization methods and bit widths. The results of GPTQ, AWQ, OmniQuant are from Shao et al. (2024), and the results of SqueezeLLM are from Kim et al. (2023). The official SqueezeLLM code does not support 2-bit quantization, and we report the available results from Kim et al. (2023).
WikiText-2 C4
Grid Method Bits 1-7B 1-13B 2-7B 2-13B 2-70B 1-7B 1-13B 2-7B 2-13B 2-70B Avg.
FP16 16 5.58 5.09 5.47 4.88 3.31 7.08 6.61 6.97 6.46 5.52 5.697
Affine GPTQ 4.00 6.13 5.40 5.83 5.13 3.58 7.43 6.84 7.37 6.70 5.67 6.008
AWQ 4.00 6.08 5.34 6.15 5.12 - 7.52 6.86 7.68 6.74 - -
OmniQuant 4.00 5.86 5.21 5.74 5.02 3.47 7.34 6.76 7.35 6.65 5.65 5.905
LeanQuantaff 4.00 5.92 5.25 5.73 5.08 3.49 7.30 6.76 7.25 6.63 5.63 5.904
Non-uniform SqueezeLLM 4.04-4.05 5.79 5.18 5.62 4.99 3.41 7.21 6.71 7.12 6.57 5.58 5.818
LeanQuantnu 4.04-4.05 5.81 5.19 5.64 4.99 3.42 7.21 6.70 7.13 6.57 5.58 5.824
Affine GPTQ 3.00 8.06 6.76 8.37 6.44 4.82 9.49 8.16 9.81 8.02 6.57 7.650
AWQ 3.00 11.88 7.45 24.00 10.45 - 13.26 9.13 23.85 13.07 - -
OmniQuant 3.00 6.49 5.68 6.58 5.58 3.92 8.19 7.32 8.65 7.44 6.06 6.591
LeanQuantaff 3.00 6.62 5.76 6.61 5.66 3.91 7.98 7.19 8.27 7.23 5.90 6.513
Non-uniform SqueezeLLM 3.02 6.32 5.60 6.18 5.36 3.77 7.75 7.08 7.72 6.97 5.83 6.258
LeanQuantnu 3.02 6.34 5.60 6.19 5.40 3.80 7.74 7.05 7.73 6.98 5.83 6.266
Affine GPTQ 2.00 1.1E5 6.8E4 3.8E4 5.6E4 2.0E4 689.13 2.5E3 NaN 323.12 48.82 NaN
OmniQuant 2.00 15.47 13.21 37.37 17.21 7.81 24.89 18.31 90.64 26.76 12.28 26.395
LeanQuantaff 2.00 18.53 14.42 25.69 24.43 7.92 19.99 16.53 27.11 20.92 10.84 18.638
Non-uniform SqueezeLLM 2.01 - N/A - 61.25 10.86 - N/A - N/A
LeanQuantnu 2.01 15.65 9.64 15.51 10.06 6.35 17.62 10.93 17.07 11.83 7.96 12.262
Table 8: Zero-shot accuracy of more quantized LLMs on benchmarks.
Method Bits ARC LAMBADA MMLU HellaS PIQA WinoG Avg.
Easy Chg Std OpenAI STEM Human. Social Other
LLaMA-7B
FP16 16 75.29 41.81 67.77 73.49 28.20 32.03 31.65 36.53 56.92 78.67 70.09 53.86
GPTQ 4.00 73.61 39.51 65.61 71.98 24.29 28.12 25.80 31.16 55.61 77.80 70.40 49.03
Affine OmniQuant 4.00 74.49 39.68 65.38 71.96 30.32 30.92 33.38 35.79 55.82 78.45 68.67 53.17
LeanQuantaff 4.00 74.03 41.64 63.75 70.56 31.27 33.13 32.66 37.79 56.31 78.40 69.30 53.53
GPTQ 3.00 66.67 35.84 49.89 58.37 25.94 27.46 23.46 24.98 51.76 75.35 64.72 43.78
OmniQuant 3.00 72.22 38.48 59.67 69.20 27.18 27.65 25.64 29.61 52.99 76.55 67.17 49.67
LeanQuantaff 3.00 73.70 40.78 65.28 72.66 27.85 31.37 30.61 35.31 56.23 78.45 70.09 52.94
GPTQ 2.00 26.35 22.01 0.00 0.00 23.69 25.08 24.28 24.01 25.69 53.70 50.99 24.95
OmniQuant 2.00 51.05 22.70 11.80 23.13 26.51 26.04 24.37 23.69 34.71 64.36 54.30 32.97
LeanQuantaff 2.00 55.98 27.22 38.56 47.45 24.45 25.31 22.85 25.46 37.19 68.12 60.77 39.40
Non-uniform SqueezeLLM 4.05 76.56 46.25 69.53 75.22 32.98 37.19 42.18 44.00 59.29 78.62 71.82 57.60
LeanQuantnu 4.05 76.39 45.05 71.55 76.48 34.76 38.77 46.12 47.18 59.29 78.78 73.24 58.87
SqueezeLLM 3.02 75.46 43.77 65.07 72.75 30.45 34.24 37.18 40.46 57.32 78.29 71.35 55.12
LeanQuantnu 3.02 75.17 43.00 70.41 77.29 33.81 38.32 43.16 45.06 57.35 78.29 71.35 57.56
SqueezeLLM 2.01 - N/A -
LeanQuantnu 2.01 50.38 24.40 31.67 41.30 21.79 24.19 21.74 24.36 37.49 65.67 58.33 36.48
LLaMA-13B
FP16 16 77.40 46.42 71.12 76.19 36.41 41.55 48.49 48.54 59.92 79.16 72.69 59.81
GPTQ 4.00 77.06 45.56 69.12 75.28 34.44 39.15 45.95 46.73 58.99 78.56 72.53 56.63
Affine OmniQuant 4.00 75.97 45.22 68.25 75.59 35.30 40.21 48.20 47.25 59.11 78.94 72.61 58.79
LeanQuantaff 4.00 76.39 46.42 70.48 76.27 35.52 39.45 46.18 47.22 58.82 78.94 72.30 58.91
GPTQ 3.00 70.92 39.93 57.29 64.82 29.37 31.94 33.18 35.34 54.05 76.99 68.43 49.13
OmniQuant 3.00 75.42 42.83 60.80 71.34 29.56 34.24 36.24 41.17 57.27 77.97 69.61 54.22
LeanQuantaff 3.00 75.84 43.34 67.49 74.85 33.37 36.56 41.92 44.74 56.75 77.97 70.48 56.66
GPTQ 2.00 27.10 21.93 0.02 0.00 23.37 25.50 23.56 24.49 25.76 53.16 49.72 24.75
OmniQuant 2.00 59.51 29.52 17.85 23.35 22.52 24.12 22.81 24.46 42.01 67.25 56.12 35.41
LeanQuantaff 2.00 61.45 29.27 44.63 50.82 28.73 26.01 27.20 27.26 39.08 71.27 65.82 42.87
Non-uniform SqueezeLLM 4.04 76.56 46.25 69.53 75.22 32.98 37.19 42.18 44.00 59.29 78.62 71.82 57.60
LeanQuantnu 4.04 76.39 45.05 71.55 76.48 34.76 38.77 46.12 47.18 59.29 78.78 73.24 58.87
SqueezeLLM 3.02 75.46 43.77 65.07 72.75 30.45 34.24 37.18 40.46 57.32 78.29 71.35 55.12
LeanQuantnu 3.02 75.17 43.00 70.41 77.29 33.81 38.32 43.16 45.06 57.35 78.29 71.35 57.56
SqueezeLLM 2.01 - N/A -
LeanQuantnu 2.01 65.66 32.42 54.49 66.93 23.44 25.50 23.98 28.42 45.66 72.80 66.14 45.95
Llama-2-13B
FP16 16 79.50 48.46 70.35 76.73 42.28 47.89 61.16 59.38 60.06 79.05 72.22 63.37
GPTQ 4.00 78.32 45.48 68.33 75.35 40.28 46.08 56.48 54.65 58.92 78.45 71.82 59.59
Affine OmniQuant 4.00 77.69 47.10 68.74 75.57 41.39 46.10 57.39 55.87 59.48 79.00 70.32 61.70
LeanQuantaff 4.00 79.42 47.27 69.16 75.90 42.21 47.31 59.90 57.93 59.07 78.24 71.82 62.57
GPTQ 3.00 72.85 39.85 59.77 67.20 34.86 38.85 47.97 46.48 54.61 76.28 70.32 53.62
OmniQuant 3.00 76.60 43.34 60.70 70.54 38.60 42.59 53.23 51.82 57.42 77.97 69.14 58.36
LeanQuantaff 3.00 77.31 44.54 68.15 75.88 37.93 43.80 53.07 52.62 56.36 76.99 70.72 59.76
GPTQ 2.00 25.84 20.22 0.00 0.00 22.84 25.59 23.53 23.98 25.97 52.07 47.75 24.19
OmniQuant 2.00 48.19 24.66 10.21 20.14 21.34 24.21 21.77 23.85 40.16 63.00 52.33 31.81
LeanQuantaff 2.00 50.88 24.32 32.70 39.57 21.50 24.38 21.90 24.40 38.01 67.19 56.91 36.52
Non-uniform SqueezeLLM 4.04 78.91 47.70 70.00 76.23 42.72 47.89 60.19 58.32 59.74 78.73 72.77 63.02
LeanQuantnu 4.04 78.91 47.56 71.12 77.43 43.51 47.44 59.54 58.83 59.58 78.62 72.06 63.15
SqueezeLLM 3.02 77.27 43.17 66.37 73.80 38.22 44.63 55.18 53.11 58.74 77.86 69.46 59.80
LeanQuantnu 3.02 77.19 44.20 71.14 78.59 40.72 45.46 56.87 55.10 56.38 77.75 70.09 61.23
SqueezeLLM 2.01 - N/A -
LeanQuantnu 2.01 62.46 30.20 47.00 61.09 25.28 27.74 27.56 28.87 42.20 69.91 62.04 44.03
Appendix G Quantization Time Cost
The time cost of LeanQuant for different models and configurations are presented in Table 9.
Table 9: Total time taken by LeanQuant for quantizing different-sized LLMs, using a single L40s-48GB GPU, an AMD EPYC 7R13 48-Core CPU, and 370GB of RAM. Llama-3.1-405B is quantized using 2 Quadro RTX 8000 GPUs, an AMD EPYC 7742 64-Core CPU, and 1.48TB of RAM.
Model Grid Group Size Bits Time
Llama-2-7B Affine - 4.00 14 mins
Affine 128 4.25 15 mins
Non-uniform - 4.05 35 mins
Llama-3-8B Affine - 4.00 16 mins
Affine 128 4.25 20 mins
Non-uniform - 4.05 37 mins
Llama-2-70B Affine - 4.00 178 mins
Non-uniform - 4.04 335 mins
Mistral-Large-Instruct-2407 (123B) Affine - 4.00 252 mins
Llama-3.1-405B Affine 128 4.25 1241 mins
Appendix H Ablation Study
Sensitivity to Hyperparameter $p$ Ablative experiments on the effects of the hyperparameter $p$ on the quality of LeanQuant models are presented in Table 10. In the case of $p=0$, the inverse Hessian diagonals are ignored as the weights for clustering, and the centroids are learned based on the density of weights. It is worth noting that $p=0$ results in sub-optimal model quality compared to higher values of $p$, which means that the loss-error-awareness property of the quantization grid is critical for maintaining model quality.
Grid Point Initialization Ablative experiments comparing k-means++ initialization with our proposed uniformly spaced grid initialization are presented in Table 11.
Table 10: The perplexity of LeanQuant models on WikiText2 and C4, using different values of $p$.
WikiText2 C4
Grid Hyperparameter 4-bit 3-bit 2-bit 4-bit 3-bit 2-bit
$p=0$ 12.13 29.92 5,991.18 16.73 22.71 5,998.96
$p=2$ 5.39 5.98 25.09 7.89 8.50 20.27
$p=3$ 5.37 5.92 22.32 7.88 8.48 19.81
Non-uniform $p=4$ 5.38 5.96 25.61 7.88 8.47 21.65
$p=0$ 14.52 80.54 230.66 16.94 69.04 243.65
$p=2$ 5.52 8.58 55.50 8.03 16.84 41.99
$p=3$ 5.51 6.36 18.33 8.03 8.80 20.20
Mistral-7B Affine $p=4$ 5.51 6.31 18.00 8.02 8.86 20.47
$p=0$ 5.69 6.76 NaN 7.15 8.23 62.00
$p=2$ 5.65 6.30 17.16 7.13 7.83 19.14
$p=3$ 5.64 6.25 17.84 7.13 7.80 19.55
Non-uniform $p=4$ 5.64 6.28 15.82 7.14 7.83 18.89
$p=0$ 5.84 8.19 93.01 7.30 9.54 85.62
$p=2$ 5.77 7.33 27.82 7.27 8.83 28.86
$p=3$ 5.75 6.80 25.97 7.26 8.32 27.57
Llama-2-7B Affine $p=4$ 5.75 6.69 26.82 7.25 8.29 28.14
Appendix I Loss Error Comparison
A comparison of the sum of loss errors $\epsilon$ between GPTQ and LeanQuant (affine and non-uniform) is presented in Figure 4.
Figure 4: Comparison of loss errors $\epsilon$ of each layer for GPTQ and LeanQuant (affine and non-uniform) during iterative quantization.
Table 11: Ablative experiments on grid point initialization.
Llama-2-7B Llama-3-8B Mistral-7B
Grid Init. 4-bit 3-bit 2-bit 4-bit 3-bit 2-bit 4-bit 3-bit 2-bit
WikiText2 K-means++ 5.64 6.25 17.84 6.59 8.31 46.31 5.37 5.92 22.32
Uniformly Spaced (ours) 5.66 6.20 17.53 6.59 7.88 41.78 5.40 5.88 19.06
C4 K-means++ 7.13 7.80 19.55 10.17 12.53 39.86 7.88 8.48 19.81
Uniformly Spaced (ours) 7.14 7.72 18.75 10.20 12.16 36.00 7.91 8.42 17.85